OpenStudy (thequestion):

Solve, using the substitution method. d + e = 2 d – e = 4 A. There are an infinite number of solutions. B. The solution is (3, –1). C. There is no solution. D. The solution is (5, 1).

OpenStudy (anonymous):

add both u will get 2d=6 d=6/2=3 now e+d=2 put d e+3=2 so substract by 3 on both sides e=-1

OpenStudy (anonymous):

answer b is correct